OpenStudy (anonymous):

solve the equation for exact solutions over the interval [0,2pi) 6 tan 3x=6

OpenStudy (anonymous):

6 tan 3x=6 tan 3x=1=tan pi/4=tan (n pi +pi/4),where n=0,1,2,3,4,5 \[3 x=\left( n \pi+\frac{ \pi }{4 } \right),x=\frac{ n \pi }{ 3 }+\frac{ \pi }{12 }\]